OpenStudy (anonymous):

PLEASE HELP! ♥ Solve the system of equations. y = 2x² – 3 y = 3x – 1

OpenStudy (anonymous):

for what? i mean slope or intersection or what

OpenStudy (anonymous):

not sure, it doesnt say

OpenStudy (zehanz):

You can do this: 2x² - 3 = 3x - 1, because both are equal to y. Now it is a quadratic equation. If you write it as 2x² - 3x - 2 = 0, you can use the Quadratic Formula to solve it for x.

OpenStudy (zehanz):

Or, maybe you can solve it by factoring.

OpenStudy (anonymous):

a. no solution b. (-1/r2, 5), (2, –5/2) c. (–1/2, –5/2), (2, 5) d. (1/2, 5/2), (2, 5)

OpenStudy (zehanz):

Do you know the Quadratic Formula?

OpenStudy (anonymous):

no i dont

OpenStudy (zehanz):

It is the formula that gives you the solutions of any quadratic equation: If \(ax^2+bx+c=0\), then \(x=\dfrac{-b \pm \sqrt{b^2-4ac}}{2a}\). So all you have to do is: look at the values of a, b and c in your equation: 2x² - 3x - 2 = 0, so a = 2 b = -3 c = -2 Plug these values in the Quadractic formula and out come your solutions! Once you've the solutions for x, you can substitute them in one of your original equations (y=3x-1 is the simplest), to calculate y. THen you have two pairs (x, y). Compare them with the answer options to see which one to choose.

OpenStudy (zehanz):

If you don't know the Quadratic Formula, chances are that up until now you always have solved quadratic equations by factoring. This is a good technique, although it only works if the solutions are "nice" numbers, like integers or fractions like ½, ¼ and so on. Looking at the y-values of the answer options, this seems the case. If we want to solve 2x² - 3x - 2 = 0 by factoring, we must write this equation as: (a + b)(c + d), which, aftwer FOILing (calculating all possible multiplications of a and b with c and d) should give 2x² - 3x - 2. This is not so hard as it seems, because although we have to come up with a, b, c and d, most of these numbers are already given! FOIL means: First, Outer, Inner, Last, so (a + b)(c + d)= (First) ac + (Outer) ad + (Inner) bc + (Last) bd all in all, this gives ac+ad+bc+bd. Now if we are to get 2x² - 3x - 2, we have to write it like this: (2x ...)(x ...), because only then 2x times x is 2x² (=the First of FOIL). Also, we don't have so much to choose for the Last part, because it means that the two numbers that go on the dots have product -2. If these numbers are "nice" as I claimed earlier, then they sould be something like 1 and -2 (product is -2), or -1 and 2 (also product -2), or even 4 and -½. Now what is it? Look at (2x+1)(x-2) = 2x² -4x+1x-2, so the middle two terms are -3x when added. So we have found the numbers we were looking for! 2x² - 3x - 2 = 0 is thus the same as (2x + 1)(x - 2) = 0. You should be able to spot the solutions for x now...